Solve the given initial-value problem. (x + y)2 dx + (2xy + x2 – 8) dy = 0, y(1) = 1 (x + y)3 (x + y)2 - 8x = -1
Answer #1

As per the given clata Let M(x, y) = (x+y) and N(x,y) = 2x + x²-8 The given clifferential equation is M(x,y) dx + (x,y) dy zowhich g(y)=-g therefore gly)=-8y + c for some 1. arbitary constante Hence u 1944) = 2/3 + 2y + xy-syze a since du = Mdx + N d